OpenStudy (anonymous):

Mr.Tanaka is 35 years old, and his son is 5. In how many years will Mr. Tanaka be three times as old as his son?

zepdrix (zepdrix):

In some unknown number of years, Mr Tanaka will be \(\large 35+x\) years old. When the same amount of years have passed, the son will be \(\large 5+x\) years old. If Mr Tanaka is 3 times as old as the son when these years have passed, we can write it like this,\[\large 35+x=3(5+x)\]From here we can solve for X, the number of years that have passed by.

OpenStudy (anonymous):

but whats the answer?

zepdrix (zepdrix):

To solve for x, first, distribute the 3 to each term inside the brackets,\[\large 35+x=15+3x\] Then subtract x from each side,\[\large 35=15+2x\]Then subtract 15 from each side,\[\large 20=2x\]Then finally divide both sides by 2. This will tell us the number of years that need to pass in order for the Father's age to be 3 times that of the son's.

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

i don't understand algebra so that's why i came on here

zepdrix (zepdrix):

That's not the purpose of this website I'm afraid - to just give you answers :( I showed you the entire process. If you're still struggling with it maybe you should get a tutor.. or study more :c Look over the steps. It's a bit of a tricky problem, I know.

OpenStudy (anonymous):

blah